# Algorithmic Fidelity of Large Language Models in Generating Synthetic German Public Opinions: A Case Study

### v1

@@ Abstract
Bolei Ma1,2    Berk Yoztyurk1    Anna-Carolina Haensch1,3    Xinpeng Wang1,2 Markus Herklotz1    Frauke Kreuter1,2,3    Barbara Plank1,2    Matthias Assenmacher1,2 1LMU Munich, 2Munich Center for Machine Learning, 3University of Maryland, College Park Equal contribution. Correspondence: bolei.ma@lmu.de
@@ 1 Introduction
Recent advances in LLMs have generated significant interest in their potential for synthetic data generation across various domains. A key and widely debated question is whether LLMs can produce synthetic data that accurately represent human opinions (Argyle et al., 2023; Santurkar et al., 2023; von der Heyde et al., 2024; Veselovsky et al., 2023; Long et al., 2024, inter alia).
Figure 1: The distribution of answer categories between November 2019 and November 2021 for the German GLES survey. The top 5 categories are highlighted. There is a significant surge in the Health Policy category from Nov. 2019 to Apr. 2020, with it becoming the dominant focus during this period and afterwards, likely due to the outbreak of COVID-19 in early 2020.
In social science research, it is usually surveys that provide insights into the attitudes and opinions of a population. Recent studies have explored using information from survey responses for LLM prompts, i.e., creating so-called personas and then using the LLM “responses” to additional questions (Argyle et al., 2023; Dominguez-Olmedo et al., 2024; Durmus et al., 2024, inter alia). Assessing the fidelity of LLMs in capturing and reproducing human opinions deepens our understanding of model behavior while at the same time helping researchers design more reliable models aligned with human values and improving their usability Ma et al. (2024). Among these studies, Argyle et al. (2023) introduced the concept of algorithmic fidelity, a framework for assessing how effectively LLMs replicate the socio-cultural context and nuanced opinions of diverse human subpopulations. Their work used LLMs to generate synthetic responses (“silicon samples”) informed by demographic and ideological profiles from political science datasets. Their findings suggest that LLMs can closely approximate real-world opinion distributions in certain contexts, such as U.S. elections, although challenges persist in fully aligning the generated data with actual demographic groups.
@@
To address these challenges, this study explores the algorithmic fidelity of LLMs in generating synthetic public opinions in an open-ended survey question based on German survey data. We use the survey question about the most important political problems in Germany from the German Longitudinal Election Study (GLES). The survey is a longitudinal panel survey, and the answer distribution can be found in Figure 1. We select variables from the original survey data to represent survey participants with their characteristics as personas. We include three LLMs (Gemma, Llama2, Mixtral) in our study and prompt them to simulate survey participants to answer the open-ended question about the most important political problem. Finally, we compare the outputs regarding the distributional alignment with original survey answers, predictive performance, and answer diversity. Our most important findings are:
1) Llama2 is better at modeling group opinions (§4.1).
2) Llama2’s representativeness fluctuates across survey waves, with the model’s representativeness of the population decreasing as survey diversity increases; the model represents subpopulation opinions unevenly, with favorable alignment for left-leaning parties (the Left, the Greens) over right parties (AfD) (§4.2).
3) Including more variables in prompts improves performance, with party affiliation being the most influential factor (§4.3).
@@ 2 Related Work
Recent studies have increasingly repurposed survey questionnaires, originally designed for public opinion polling, to assess the opinions generated by LLMs (Ma et al., 2024). For instance, Santurkar et al. (2023) identified significant differences between LLM opinion distributions and US-based survey participants. Similarly, Dominguez-Olmedo et al. (2024) highlighted disparities between LLM and human opinions, emphasizing the sensitivity of model outputs to biases in prompting. Tjuatja et al. (2024) found that LLMs are highly sensitive to prompt perturbations and fail to replicate human-like behavior. Collectively, these studies suggest that LLMs tend to align more closely with populations holding left-leaning, Western-oriented values.
Opinion Generation in German Contexts.
@@
Previous studiesprimarily focused on closed-ended multiple-choice questions, often relying on the model’s first token prediction (e.g., Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Tjuatja et al., 2024) or semi-automated extraction of text answers (von der Heyde et al., 2024). Alternatively, Wang et al. (2024a, b) proposed training a classifier directly on LLM responses manually labeled by annotators, finding this method more robust. After the output extraction, a few evaluation metrics have been applied to measure the alignment of human and LLM responses Ma et al. (2024), such as Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023), 1-Wasserstein distance Santurkar et al. (2023); Hwang et al. (2023), KL divergence Dominguez-Olmedo et al. (2024); Sun et al. (2024), Euclidean distance Wang et al. (2023), Jensen-Shannon distance Durmus et al. (2024), etc. and correlation and statistical analyses Sun et al. (2024); Jiang et al. (2024b). For our case study, we adapt these metrics to examine the fidelity of LLM-generated synthetic German public opinions.
@@ 3.1 Data
We use the GLES Panel dataset from GESIS (2023). The survey consists of 21 waves [^2] A survey wave refers to a single round of data collection in a panel survey, gathering information from the same participants at multiple points in time Andreß et al. (2013). and contains socio-demographic information, vote intentions, choices, and political attitudes of participants. The target population is German citizens eligible to vote during the respective elections in Germany. Along with the respondents’ socio-demographic data for our prompts, we use respondents’ answers to the question “What is the most important (political) problem in Germany?” for the comparison of human answers and LLM outputs. The answers of participants were collected as free-form texts.
@@ 3.2 Text Generation
We chose three instruction-tuned open-weight LLMs: Llama-2-13b-chat-hf Touvron et al. (2023), Gemma-7b-it Team et al. (2024), and Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024a).
@@ 4.2 Experiment 2: Wave Experiment with Llama2
A case study on party variables: Llama2 better models groups with left-leaning parties. To check how much of the JS distance can be associated with the modeling difficulty of the variables, we plot the subpopulation entropy and JS distances for leaning party in Figure 5. Since the population entropy is the same for all subpopulations, we can safely assume that lower conditional entropy means higher information gain for that variable. This allows us to examine how representativeness, linked to the available information in a variable, impacts alignment success. When mutual information is high, LLMs can better model subpopulation behavior. However, certain groups do not fall into this trend line; “die Linke (the Left)” and “die Grünen (the Greens)” are modeled better, and “AfD” is worse (compared to the information their groups carry). This finding aligns with previous work (Santurkar et al., 2023; von der Heyde et al., 2024, e.g.,), which shows that LLMs tend to have a more left-leaning feature. We present additional results on other variables and find similar results as previous work showing that LLMs are biased towards left, Western, and educated people in §G.
@@ 4.3 Experiment 3: Ablation Experiment
  • Using only one social group variable (1-var-*),
  • Excluding one social group variable (w/o-*),
  • Using no social group variables (LLM-base).
These were compared against the experiment with all variables included (LLM-all).
@@
LLM outputs show stronger correlation with variables when prompted with only one variable. Table 3 compares Cramér’s V values between the survey, 1-var, and Llama2-all experiments. When only one variable is provided, the generated texts show stronger correlations with the input variable. Although JS distances decrease when more variables are added, this observation suggests that synthetic data patterns are dynamic and can be influenced by the number of prompt variables.
@@
Figure 7: Information Gain for leaning party variable, comparing survey entropy to Llama2-all, 1-var-party, and Llama2-base. Left: population entropy ($H(Y)$), right: subpopulation entropy ($H(Y\mid X_{i})$).
The inclusion of the party variable has the most significant impact on model performance, with its presence leading to substantial improvements in information gain. As shown in Figure 6, adding either only the party (1-var-party) or education variable alone results in the greatest reduction in JS distance compared to the model without demographics; excluding only the party variable leads to a smaller improvement in JS distance, both highlighting the party variable’s greater impact. Similar to Figure 3, we plot information gain for party, comparing survey entropy to Llama2-all, 1-var-party, and Llama2-base in Figure 7. As expected, Llama2-base, which includes no subpopulation variables, produces entropies close to the population entropy, with random variations of ±$\pm 0.03$. However, in the 1-var-party experiment, information gain ranges from 0.2 to 1.3, significantly above random variation. This, along with Cramér’s V values, suggests that the model generates typical responses, reducing the variation in subpopulation opinions. Further detailed experimental results are provided in §G of the Appendix.
@@ 5 Discussion
Modeling the German population. von der Heyde et al. (2024) found that GPT-3.5 vote predictions for the 2017 German Federal Election are inaccurate and biased towards the Green and Left parties. We also found that the Llama2 model is better at modeling the opinions of the Green and the Left parties. The subpopulation entropy and in-group diversity can partially explain this finding. Other factors could be the models’ training data and the RLHF methods used.
Reduction in in-group diversity. Bisbee et al. (2024) found that while GPT-3.5 could replicate survey averages, its synthetic answers lacked variation compared to real survey data. Similarly, von der Heyde et al. (2024) noted GPT-3.5’s difficulty in capturing nuanced subpopulation behaviors. In our analysis, we also observe a reduction in in-group diversity under certain conditions, particularly when only one variable is provided to the model (1-var-prompts) or when using the Mixtral model. This suggests that the ability to represent within-group diversity is limited by the model’s input structure and specific architecture.
The role of LLM prompts. Binz and Schulz (2023) highlighted how cognitive biases, such as the framing effect, influenced GPT-3’s outputs. In our third study, we also noticed that providing only one variable in a prompt (1-var) caused Llama2 to focus disproportionately on that variable, possibly interpreting it as more critical than when multiple variables were included. Interestingly, this effect varied by model: Mixtral over-relied on variables even with full prompts, while Llama2 showed stronger biases with fewer variables. This suggests that model-specific differences influence how demographic factors are integrated and highlights the need for careful prompt design. Argyle et al. (2023) motivated the silicon sampling approach on this conditional probability formula:
@@ 6 Conclusion and Recommendations
This paper evaluates the ability of LLMs to represent the opinions of German subpopulations. While von der Heyde et al. (2024) identified challenges with GPT-3.5 in capturing the nuances of German subpopulations and the unique dynamics of Germany’s multi-party system in a closed-ended setup focused on voting choice, our approach shifts to free-form text responses. We explore how these responses align with survey data, focusing on more open-ended questions to better reflect the complexities of public opinion. By using free-form text responses rather than multiple-choice questions, we can identify detailed issues in contextual information and the variety of different subpopulations, underscoring the value of this evaluation approach. Our findings show that LLMs, particularly the Llama2 model, are capable of associating text responses with social-demographic variables, indicating a degree of representativeness. However, the number of variables included in the prompt plays a crucial role in model performance. Despite this, the models still tend to generate stereotypical representations, with a noticeable favor towards left-leaning parties, consistent with previous findings on the limited diversity of opinions reflected in LLMs.
@@
Additionally, our study focused on three open-weight LLMs and the German language, which could limit the scope of our findings and might have restricted the generalizability of our findings to other LLMs or languages. Future research should include a broader variety of models and prompting languages to explore their performance and generalizability.
@@ Acknowledgments
We thank the members of SODA Lab and MaiNLP labs from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. Xinpeng Wang and Barbara Plank were funded by ERC Consolidator Grant DIALECT (101043235). Markus Herklotz and Matthias Assenmacher were funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 27/1 - 460037581.

### v2

@@ Abstract
Bolei Ma1,2    Berk Yoztyurk1    Anna-Carolina Haensch1,2,3    Xinpeng Wang1,2 Markus Herklotz1    Frauke Kreuter1,2,3    Barbara Plank1,2    Matthias Assenmacher1,2 1LMU Munich, 2Munich Center for Machine Learning, 3University of Maryland, College Park Equal contribution. Correspondence: bolei.ma@lmu.de
@@ 1 Introduction
Recent advances in LLMs have generated significant interest in their potential for synthetic data generation across various domains. A key and widely debated question is whether LLMs can produce synthetic data that accurately represent human opinions (Argyle et al., 2023; Santurkar et al., 2023; Veselovsky et al., 2023; von der Heyde et al., 2025; Long et al., 2024, inter alia).
Figure 1: The distribution of the top 5 answer categories between November 2019 and November 2021 in the German GLES survey. There is a significant surge in the Health Policy category from November 2019 to April 2020, with it becoming the dominant focus during this period and afterwards, likely due to the outbreak of COVID-19 in early 2020.
In social science research, it is usually surveys that provide insights into the attitudes and opinions of a population. Recent studies have explored using information from survey responses for LLM prompts, i.e., creating so-called personas and then using the LLM “responses” to additional questions (Argyle et al., 2023; Dominguez-Olmedo et al., 2024; Durmus et al., 2024, inter alia). Assessing the fidelity of LLMs in capturing and reproducing human opinions deepens our understanding of model behavior while at the same time helping researchers design more reliable models aligned with human values and improving their usability Ma et al. (2024). Among these studies, Argyle et al. (2023) introduced the algorithmic fidelity, a concept for assessing how effectively LLMs replicate the socio-cultural context and nuanced opinions of diverse human subpopulations. Their work used LLMs to generate synthetic responses (“silicon samples”) informed by demographic and ideological profiles from political science datasets. Their findings suggest that LLMs can closely approximate real-world opinion distributions in certain contexts, such as U.S. elections, although challenges persist in fully aligning the generated data with actual demographic groups.
@@
To address these challenges, this study explores the algorithmic fidelity of LLMs in generating synthetic public opinions in an open-ended survey question based on German survey data. We use the survey question about the most important problems facing Germany today from the German Longitudinal Election Study (GLES). The survey is a longitudinal panel survey, and the answer distribution can be found in Figure 1. We select variables from the original survey data to represent survey participants with their characteristics as personas. We include three LLMs (Gemma, Llama2, Mixtral) in our study and prompt them to simulate survey participants to answer the open-ended question about the most important political problem in Germany. Finally, we compare the outputs regarding the distributional alignment with original survey answers, predictive performance, and answer diversity. Our most important findings are:
  • (1)
    Llama2 is better at modeling group opinions (§4.1).
  • (2)
    Llama2’s representativeness fluctuates across survey waves, with the model’s representativeness of the population decreasing as survey diversity increases; the model represents subpopulation opinions unevenly, with favorable alignment for left-leaning parties (such as the Left, the Greens) over right-parties (such as AfD) (§4.2).
  • (3)
    Including more variables in prompts improves performance, with party affiliation being the most influential factor (§4.3).
@@ 2 Related Work
Recent studies have increasingly repurposed survey questionnaires, originally designed for public opinion polling, to assess the opinions generated by LLMs (Ma et al., 2024). For instance, Santurkar et al. (2023) identified significant differences between LLM opinion distributions and US-based survey participants. Similarly, Dominguez-Olmedo et al. (2024) highlighted disparities between LLM and human opinions, emphasizing the sensitivity of model outputs to biases in prompting. Tjuatja et al. (2024) found that LLMs are highly sensitive to prompt perturbations and fail to replicate human-like behavior. Collectively, these studies suggest that LLMs align more closely with populations holding left-leaning, Western-oriented values.
Opinion Generation in the German Contexts.
@@
Previous studiesprimarily focused on closed-ended multiple-choice questions, often relying on the model’s first token prediction (e.g., Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Tjuatja et al., 2024) or semi-automated extraction of text answers (von der Heyde et al., 2025). Alternatively, Wang et al. (2024a, b) proposed training a classifier directly on LLM responses manually labeled by annotators, finding this method more robust. After the output extraction, a few evaluation metrics have been applied to measure the alignment of human and LLM responses Ma et al. (2024), such as Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023), 1-Wasserstein distance Santurkar et al. (2023); Hwang et al. (2023), KL divergence Dominguez-Olmedo et al. (2024); Sun et al. (2024), Euclidean distance Wang et al. (2023), Jensen-Shannon distance Durmus et al. (2024), etc. and correlation and statistical analyses Sun et al. (2024); Jiang et al. (2024b). For our case study, we adapt these metrics to examine the fidelity of LLM-generated synthetic German public opinions.
@@ 3.1 Data
We use the GLES Panel dataset from GESIS (2023). The survey consists of 21 waves [^2] A survey wave refers to a single round of data collection in a panel survey, gathering information from the same participants at multiple points in time Andreß et al. (2013). and contains socio-demographic information, vote intentions, choices, and political attitudes of participants. The target population is German citizens eligible to vote during the respective elections in Germany. Along with the respondents’ socio-demographic data for our prompts, we use respondents’ answers to the question “In your opinion, what is the most important problem facing Germany today?” for comparing human answers and LLM outputs. The answers of participants were collected as free-form texts.
@@ 3.2 Text Generation
We chose three instruction-tuned open-weight LLMs: Llama-2-13b-chat-hf Touvron et al. (2023), Gemma-7b-it Team (2024), and Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024a).
@@ 4.2 Experiment 2: Wave Experiment with Llama2
A case study on party variables: Llama2 better models groups with left-leaning parties. To check how much of the JS distance can be associated with the modeling difficulty of the variables, we plot the subpopulation entropy and JS distances for leaning party in Figure 5. Since the population entropy is the same for all subpopulations, we can safely assume that lower conditional entropy means higher information gain for that variable. This allows us to examine how representativeness, linked to the available information in a variable, impacts alignment success. When mutual information is high, LLMs can better model subpopulation behavior. However, certain groups do not fall into this trend line; “die Linke (the Left)” and “die Grünen (the Greens)” are modeled better, and “AfD” is worse (compared to the information their groups carry). This finding aligns with previous work (Santurkar et al., 2023; von der Heyde et al., 2025, e.g.,), which shows that LLMs tend to have a more left-leaning feature. We show additional results on other variables and observe similar results showing that LLMs are biased towards left, Western, and educated people in §G (4th paragraph).
@@ 4.3 Experiment 3: Ablation Experiment
  • Including only one social group variable.
  • Excluding one social group variable.
  • Using no social group variables.
These were compared against the experiment with all variables included.
@@
LLM outputs show stronger correlation with variables when prompted with only one variable. Table 3 compares Cramér’s V values between the survey, including only one variable, and all variables included experiments. When only one variable is provided, the generated texts show stronger correlations with the input variable. Although JS distances decrease when more variables are added, this observation suggests that synthetic data patterns are dynamic and can be influenced by the number of prompt variables.
@@
Figure 7: Information Gain for leaning party variable, comparing survey entropy to Llama2-all (with all variables), 1VAR-party (with only party variable), and Llama2-base (with no variables). Left: population entropy ($H(Y)$), right: subpopulation entropy ($H(Y\mid X_{i})$).
The inclusion of the party variable has the most significant impact on model performance, with its presence leading to substantial improvements in information gain. As shown in Figure 6, adding either only the party or education variable alone results in the greatest reduction in JS distance compared to the model without demographics; excluding only the party variable leads to a smaller improvement in JS distance, both highlighting the party variable’s greater impact. Similar to Figure 3, we plot information gain for party, comparing survey entropy to Llama2 with all variables, with only one party variable, and with no variables in Figure 7. As expected, Llama2-base, which includes no subpopulation variables, produces entropies close to the population entropy, with random variations of ±$\pm 0.03$. However, in the 1VAR-party experiment, information gain ranges from 0.2 to 1.3, significantly above random variation. This, along with Cramér’s V values, suggests that the model generates typical responses, reducing the variation in subpopulation opinions. Further detailed experimental results are provided in §G of the Appendix.
@@ 5 Discussion
Algorithmic Fidelity in Modeling the German population. von der Heyde et al. (2025) found that GPT-3.5 vote predictions for the 2017 German election are inaccurate and biased towards the Green and Left parties. We also found that the model is better at modeling the opinions of the Green and the Left parties than the right parties. The subpopulation entropy and in-group diversity can partially explain this finding. Other factors could be the models’ training data and the RLHF methods used.
Reduction in in-group diversity. Bisbee et al. (2024) found that while GPT-3.5 could replicate survey averages, its synthetic answers lacked variation compared to real survey data. Similarly, von der Heyde et al. (2025) noted GPT-3.5’s difficulty in capturing nuanced subpopulation behaviors. In our analysis, we also observe a reduction in in-group diversity under certain conditions, particularly when only one variable is provided to the model or when using the Mixtral model. This suggests that the ability to represent within-group diversity is limited by the model’s input structure and specific architecture.
The role of LLM prompts. Binz and Schulz (2023) highlighted how cognitive biases, such as the framing effect, influenced GPT-3’s outputs. In our third study, we also noticed that providing only one variable in a prompt caused Llama2 to focus disproportionately on that variable, possibly interpreting it as more critical than when multiple variables were included. Interestingly, this effect varied by model: Mixtral over-relied on variables even with full prompts, while Llama2 showed stronger biases with fewer variables. This suggests that model-specific differences influence how demographic factors are integrated and highlights the need for careful prompt design. Argyle et al. (2023) motivated the silicon sampling approach on this conditional probability formula:
@@ 6 Conclusion and Recommendations
This paper evaluates the algorithmic fidelity of LLMs to represent the opinions of German subpopulations. While von der Heyde et al. (2025) found that GPT-3.5 struggles with the nuances of German subpopulations and the country’s multi-party system in closed-ended voting questions, we explore instead free-form open-ended text responses, focusing on how these responses align with survey data. By using free-form text responses rather than multiple-choice questions, we can identify detailed issues in contextual information and the variety of different subpopulations, underscoring the value of this evaluation approach. Our findings show that LLMs, particularly the Llama2 model, are capable of associating text responses with social-demographic variables, indicating a degree of representativeness. However, the number of variables included in the prompt plays a crucial role in model performance. Despite this, the models still tend to generate stereotypical representations, with a noticeable favor towards left-leaning parties, consistent with previous findings on the limited diversity of opinions reflected in LLMs.
@@
Additionally, our study focused on three open-weight LLMs and the German language, which could limit the scope of our findings and might have restricted the generalizability of our findings to other LLMs or languages. Future research should include a broader variety of models and prompting languages to explore their performance and generalizability. Moreover, as our approach primarily relies on zero-shot prompting, a promising next step would be to evaluate how alternative methods, such as fine-tuning (Cao et al., 2025; Holtdirk et al., 2025, e.g.,), could influence model performance in the German context.
@@ Acknowledgments
We thank the members of SODA Lab and MaiNLP labs from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. Xinpeng Wang and Barbara Plank are supported by ERC Consolidator Grant DIALECT (101043235). Markus Herklotz and Matthias Assenmacher are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 27/1 - 460037581.